OpenStudy (anonymous):

What is the minimum speed a ball must be thrown upward if it is to reach a height of 11 feet?

OpenStudy (anonymous):

What part has you lost. Do you know the basic equations of motion for projectiles?

OpenStudy (anonymous):

honestly I don't. This is my calc hw and I have never taken physics.

OpenStudy (anonymous):

Ah okay. Well let's call the position x(t), velocity v(t)=x'(t) and acceleration a(t)=x''(t) We know that x''(t)=-g. So you can play around with this. One approach would be to find the time when x'(t)=0 which is the time of maximum height. Then see if you can't get x(t_max) to evaluate to 11

OpenStudy (anonymous):

-.5gt^2 = x(t) correct?

OpenStudy (anonymous):

You are forgetting the initial consitiond -- the constants of integration shouldnt just disappear. Commonplace is to do x(0)=x0 and x'(0)=v0

OpenStudy (anonymous):

ya i am lost. sorry

OpenStudy (anonymous):

I mean how did you get from x''(t)=-g to your x(t)?

OpenStudy (anonymous):

the integral of -g = -gt+c and then the integral of -gt+c = -.5gt^2+ct+C. I assuming the c's are zero because i am starting at the origin

OpenStudy (anonymous):

Well one of the cs will not be. When you go from x''(t)=-g to x'(t)=-gt+c you need to set x'(0) to the initial condition which we will just call v0 for now. So x'(0)=v0. So that c can be solved for. Generally you do the same thing for position by setting x(0)=x0

OpenStudy (anonymous):

i will give that a shot

OpenStudy (anonymous):

after attempting your approach i get a solution of 8sqrt(11). that look correct?

OpenStudy (anonymous):

sorry all of our solutions have to be in square roots so some sort.

OpenStudy (anonymous):

What did you get for x(t)?

OpenStudy (anonymous):

after analysis the question a little further i realized that i only need to find the velocity so i asked a buddy and he gave me the equation of vf^2=v0^2+2ad which i transposed to sqrt(Vf-2AD) which came out to sqrt(0-2*32*11)

OpenStudy (anonymous):

Well sure the physics approach is easier knowing the equations but I thought you wanted to do it with calc.

OpenStudy (anonymous):

i do lol, my friend said this is easier approach and I was lost with the calc approach. this is just the 2 week of calc and it has been crazy since it is online.